OpenStudy (anonymous):

x^2-4x+___=(__)^2 what numbers would make this true ?

OpenStudy (anonymous):

You need to do whats called completing the square. when you have a quadratic equation in the form: \[x^2+bx+c\] for it to be a square, c must be: \[c = (\frac{b}{2})^2\]

OpenStudy (anonymous):

+4; this would give (x-2)^2

OpenStudy (anonymous):

Would the 2 be negative or positive ?

OpenStudy (anonymous):

negative. (x-2)(x-2)=x^2-4x+4

OpenStudy (anonymous):

- Thanks , alot .

OpenStudy (anonymous):

no prob :)